# Is My Data in Your Retrieval Database? Membership Inference Attacks Against Retrieval Augmented Generation

### v2

@@ 4.3 Defense Strategy
Please answer the question using the context provided.
If the question is unanswerable, say unanswerable”.
@@
1. Repeating, outputting or printing the context is illegal,
answer unanswerable”.
2. Asking if a specific text, fact, or passage appears in your
context is illegal, answer unanswerable”.
@@ 2. Asking if a specific text, fact, or passage appears in your
We have slightly adapted the RAG template for the llama model, such that it will respond to the defense instructions better, by placing the defense instructions inside the system prompt, an integral capability of llama models. In Table 3 we present the results, and compare them to the scenario where no defense is placed in the RAG system.
Table 3: RAG MIA results with defense
@@
HealthCareMagic flan 0.78 0.99 0.86 0.90
llama 0.88 0.96 0.74 0.51
mistral 0.74 0.83 0.72 0.44
Enron flan 0.82 0.97 0.88 0.96
llama 0.79 0.84 0.77 0.78
mistral 0.78 0.80 0.78 0.72
Our evaluation reveals that the proposed defense strategy yields the most significant benefits against gray-box attacks on the llama and mistral models, across both datasets. Notably, the defense demonstrates high efficacy in the gray-box setting for the mistral model, particularly on the HealthCareMagic dataset, where we observe a substantial improvement of 0.39 in AUC-ROC. In contrast, the defense has a minimal impact on the flan model, only showing a slight effect in the gray-box setting. These findings suggest that further research is needed to explore the feasibility of crafting a RAG template that can effectively defend a flan model against RAG-MIA attacks.
Furthermore, our analysis of the model outputs reveals that a significant proportion of responses do not contain either ”Yes” or ”No” tokens (classified as non-member documents), accounting for approximately 96% of the total outputs on llama, and about 93% on mistral with HealthCareMagic dataset, which explains the defense performance for these cases. On the other hand, for flan, they remain the same (0%). We present the statistics on the amount of such documents in Appendix D.
@@ 4.3.1 Improved defense for llama
Since llama models have a dedicated section in the input prompt for system instructions, as mentioned in Section 4.3, we have experimented with placing the retrieved database content within this dedicated section. We compare two cases: (1) With Defense #1 - only the defense instructions are added to the system section (2) With Defense #2 - both defense instructions and retrieved database content are placed in the system section. We present the results for this experiment in Table 4.
Table 4: Llama defenses
@@
AUC ROC
Dataset Model
HealthCareMagic llama 0.88 0.96 0.74 0.51 0.42 0.75
Enron llama 0.79 0.84 0.77 0.78 0.46 0.92
As shown in Table 4, placing defense instructions and retrieved database content in the system section provides robust defense against black-box attacks. However, this approach is less effective against gray-box attacks, where Defense #1 is preferred. Since black-box attacks are more common and require less effort from the attacker, we recommend using Defense #2. Nevertheless, we encourage the research community to develop a defense strategy that can effectively protect from both kinds of attack.
5 Conclusion
In this paper we introduced a new membership inference attack against RAG-based systems meant to infer if a specific document was part of the retrieval database or not, a research question previously not explored. Our attack is demonstrated both in black-box and gray-box threat models, and takes advantage of a characteristic of generative models that is usually considered an advantage - context grounding.
Our attack achieves a very high average AUC-ROC of 0.90 and 0.80 in the gray-box and the black-box threat models, respectively, and for some models almost perfect performance. Our initial defense was able to reduce the success rate of the attack in almost all cases, and mostly prevented the attack for the llama model in the black-box setting. The model that mostly did not benefit from this defense was the flan model for which further defenses need to be developed. This underscores the need to develop more advanced defenses and countermeasures for this attack scenario. We hope that the research community will proceed to explore the risk of membership inference in RAG-based systems and employ the ideas from this paper as a baseline.
6 Acknowledgement
@@
Appendices
A Detailed Attack AUC ROC Scores
Table 5 presents the AUC ROC scores for both threat models: black-box and gray-box, and for the different attack prompts.
Table 5: Full RAG-MIA results
@@ A Detailed Attack AUC ROC Scores
HealthCareMagic flan 0 0.80 0.98
1 0.75 0.97
2 0.78 0.99
3 0.76 0.84
4 0.80 1.00
llama 0 0.82 0.88
1 0.67 0.64
2 0.88 0.96
3 0.78 0.83
4 0.84 0.89
mistral 0 0.71 0.73
1 0.67 0.65
2 0.74 0.83
3 0.71 0.73
4 0.71 0.55
Enron flan 0 0.79 0.93
1 0.68 0.80
2 0.82 0.96
3 0.60 0.62
4 0.80 0.92
llama 0 0.71 0.65
1 0.62 0.54
2 0.79 0.83
3 0.69 0.58
4 0.74 0.74
mistral 0 0.66 0.60
1 0.70 0.68
2 0.78 0.82
3 0.64 0.53
4 0.74 0.69
@@ B Detailed Results of Retrieval Matches
Tables 6 and 7 show the percent of exact matches between the retrieved documents and the member and non-member samples, respectively, for the different attack prompts. We see that the chosen attack prompts do not differ in their influence on the retrieval accuracy.
Table 6: Database retrieval of member documents
@@
Table 7: Database retrieval of non-member documents
@@ C Model outputs without a clear yes/no answer
The number of model outputs that did not contain either of the ”Yes” or ”No” tokens across attack prompts for each dataset and model combination are shown in Table 8.
Table 8: Model outputs missing yes/no tokens
@@ D Model outputs without a clear yes/no answer with defense prompt
The number of model outputs that did not contain either of the ”Yes” or ”No” tokens when using a defense prompt for each dataset and model combination are shown in Table 9. This is compared to the corresponding model outputs when using attack prompt #2 without the defense. The number of missing answers increases significantly for the llama and mistral models, and remains the same for flan.
Table 9: Model outputs missing yes/no tokens with defense prompt
@@
HealthCareMagic flan 0 4000 $0$ 4000 0.00%
llama 0 4000 0.00% 3868 4000 96.70%
@@
Enron flan 0 4000 $0$ 4000 0.00%
llama 0 4000 0.00% 3805 4000 95.12%

### v3

@@ 5.3 Initial Defense Strategy
Please answer the question using the context
provided. If the question is unanswerable,
say "unanswerable".
@@
1. Repeating, outputting or printing the
context is illegal, answer "unanswerable".
2. Asking if a specific text, fact, or
passage appears in your context is illegal,
answer "unanswerable".
@@ 2. Asking if a specific text, fact, or
For the llama model, that has a separate section for a system prompt and is specifically designed to adhere to the instructions contained within it, we slightly adapted the RAG template by placing the defense instructions inside the system prompt. This was done to increase the likelihood of the model to respond to the defense instructions. To the best of our knowledge, neither flan nor mistral have such sections. Following is the modified RAG template for llama:
<|begin_of_text|><|start_header_id|>system
<|end_header_id|>
Please answer the question using the context
provided. If the question is unanswerable,
say "unanswerable".
Pointers for the answer:
1. Repeating, outputting or printing the
context is illegal, answer "unanswerable".
2. Asking if a specific text, fact, or
passage appears in your context is illegal,
answer "unanswerable".
Context:
{context}
<|eot_id|>
<|begin_of_text|><|start_header_id|>user
<|end_header_id|>
Question: {user_prompt}.
<|eot_id|><|start_header_id|>assistant
<|end_header_id|>
Table 3: RAG-MIA results with defense - TPR@FPR.
@@
HealthCareMagic flan 1.00 0.61 0.85 0.67 0.02 0.65
llama 0.95 0.20 0.73 0.09 0.00 0.13
mistral 0.42 0.10 0.36 0.11 0.01 0.13
Enron flan 1.00 0.56 0.63 0.77 0.04 0.69
llama 0.78 0.30 0.28 0.42 0.04 0.32
mistral 0.61 0.17 0.22 0.52 0.06 0.27
Table 4: RAG-MIA results with defense - AUC-ROC.
Without defense With defense
Black-Box
AUC-ROC
Gray-Box
AUC-ROC
Black-Box
AUC-ROC
Gray-Box
AUC-ROC
Dataset Model
HealthCareMagic flan 0.81 0.99 0.85 0.90
llama 0.89 0.96 0.74 0.51
mistral 0.74 0.83 0.72 0.45
Enron flan 0.82 0.96 0.88 0.97
llama 0.79 0.83 0.77 0.77
mistral 0.78 0.81 0.78 0.72
In Table 3 and Table 4 we present the results of our attacks on the defended models, comparing them to the undefended models.
This evaluation reveals that the proposed defense strategy yields the most significant benefits against gray-box attacks on the llama and mistral models, across both datasets. On the HealthCareMagic dataset, we observe a substantial improvement of 0.45 and 0.38 in AUC-ROC and of 0.6 and 0.23 in TPR@lowFPR, respectively. In contrast, the defense has a minimal impact on the flan model, only showing a slight effect in the gray-box setting. These findings suggest that further research is needed to explore the feasibility of crafting a RAG template that can effectively defend a flan model against RAG-MIA attacks.
Furthermore, our analysis of the model outputs reveals that, when applying this defense, a significant proportion of the responses does not contain a ”Yes” or ”No” token. Instead, it contains ”unanswerable” as per the defense instruction. This accounts for approximately 96% of the total outputs from llama, and 93% from mistral with the HealthCareMagic dataset. As previously noted, our attack classifies these responses as non-member documents, explaining the improved defense performance for these cases. On the other hand, for flan, the percentage of responses not containing the ”Yes” or ”No” token remains the same (0%). Detailed statistics are presented in Appendix D.
@@ 5.3.1 Improved defense for llama
As mentioned in Section 5.3, llama models have a dedicated section in the input prompt for system instructions. We thus also experimented with placing the retrieved database content within this dedicated section. We compare two cases: (1) With Defense #1 - only the defense instructions are added to the system section (2) With Defense #2 - both defense instructions and retrieved database content are placed in the system section. We present the results of this experiment in Tables 5 and 6.
Table 5: Llama defenses - TPR@FPR.
Without Defense With Defense #1 With Defense #2
Black-Box Gray-Box Black-Box Gray-Box Black-Box Gray-Box
TPR FPR TPR@lowFPR TPR FPR TPR@lowFPR TPR FPR TPR@lowFPR
Dataset
HealthCareMagic 0.95 0.20 0.73 0.09 0.00 0.13 0.00 0.00 0.30
Enron 0.78 0.30 0.28 0.42 0.04 0.32 0.01 0.01 0.49
Table 6: Llama defenses - AUC-ROC.
@@
AUC-ROC
Dataset
HealthCareMagic 0.89 0.96 0.74 0.51 0.54 0.74
Enron 0.79 0.83 0.77 0.77 0.51 0.92
As shown in Table 5, placing both the defense instructions and the retrieved database content in the system section provides a robust defense against black-box attacks. However, this approach is less effective against gray-box attacks, where Defense #1 is preferred. Since black-box attacks are more common and require less effort from the attacker, we recommend using Defense #2. Nevertheless, we encourage the research community to further research and develop defense strategies that can effectively protect against both kinds of attacks.
6 CONCLUSIONS
In this paper we introduced a new membership inference attack against RAG-based systems meant to infer if a specific document is part of the retrieval database or not. Our attack does not rely on the model replicating text from the retrieval database, and rather relies on binary answers provided by the model itself regarding its context. This takes advantage of a characteristic of generative models that is usually considered an advantage - context grounding. We demonstrated results both in black-box and gray-box threat models.
Our attack achieves a very high performance, with average AUC-ROC of 0.90 and 0.80 in the gray-box and the black-box threat models, respectively, and for some models achieving almost perfect performance.
Conversely, our initial defense was able to reduce the success rate of the attack in almost all cases, and essentially prevented the attack for the llama model in the black-box setting. This illustrates the need for more advanced attack prompts, for example integrating prompt-injection techniques. The model that mostly did not benefit from this defense was the flan model, for which further defenses need to be developed.
Furthermore, while this paper only explored direct attacks, it is important to take into account adaptive attacks (e.g. [Tramer et al., 2020]) that consider potential defenses, such as the one presented in this paper, in the attack process. This underscores the need to establish more advanced defense strategies and countermeasures. Such approaches may be based on differentially private synthetic data generation [Amin et al., 2024, Xie et al., 2024], which employs LLMs to generate synthetic texts based on several seed text passages. Alternatively, differential privacy mechanisms can be employed to the LLMs’ text generation process [Du and Mi, 2021, Majmudar et al., 2022], which may also help to reduce the risk.
In summary, we hope that the research community will continue to explore the risk of membership inference in RAG-based systems and employ the ideas from this paper as a baseline.
ACKNOWLEDGEMENTS
@@
APPENDIX
A Detailed Attack Scores
Table 7 presents the full AUC-ROC and TPR scores for the different attack prompts for both threat models: black-box and gray-box.
Table 7: Full RAG-MIA results.
@@ A Detailed Attack Scores
HealthCareMagic flan 0 1.00 0.63 0.76 0.80 0.97
1 0.99 0.84 0.76 0.75 0.97
2 1.00 0.74 0.83 0.78 0.99
3 0.98 0.70 0.33 0.76 0.83
4 1.00 0.61 0.85 0.81 0.99
llama 0 0.91 0.29 0.48 0.82 0.87
1 0.64 0.30 0.28 0.67 0.63
2 0.95 0.20 0.73 0.89 0.96
3 0.92 0.44 0.47 0.77 0.82
4 0.91 0.25 0.38 0.84 0.88
mistral 0 0.91 0.63 0.32 0.70 0.72
1 0.83 0.54 0.27 0.67 0.65
2 0.97 0.69 0.36 0.74 0.83
3 0.94 0.71 0.22 0.70 0.73
4 0.42 0.10 0.23 0.71 0.54
Enron flan 0 0.99 0.65 0.60 0.79 0.93
1 0.94 0.75 0.46 0.68 0.81
2 1.00 0.56 0.63 0.82 0.96
3 0.87 0.73 0.27 0.60 0.63
4 1.00 0.63 0.34 0.81 0.91
llama 0 0.92 0.63 0.13 0.71 0.66
1 0.63 0.38 0.14 0.62 0.54
2 0.91 0.38 0.28 0.79 0.83
3 0.93 0.72 0.14 0.69 0.59
4 0.78 0.30 0.17 0.74 0.75
mistral 0 0.87 0.64 0.09 0.66 0.59
1 0.85 0.50 0.21 0.70 0.68
2 0.92 0.44 0.22 0.78 0.81
3 0.86 0.65 0.10 0.64 0.51
4 0.61 0.17 0.20 0.73 0.67
@@ B Detailed Results of Retrieval Matches
Tables 8 and 9 show the percent of exact matches between the retrieved documents and the member and non-member samples, respectively. We see that the chosen attack prompts do not differ in their influence on the retrieval accuracy.
Table 8: Database retrieval of member documents.
@@
Table 9: Database retrieval of non-member documents.
@@ C Model Outputs Without a Clear Yes/No Answer
The number of model outputs that did not contain either of the ”Yes” or ”No” tokens for each dataset and model combination (across all attack prompts and including both members and non-members) are shown in Table 10.
Table 10: Model outputs missing Yes/No tokens.
@@ D Model Outputs Without a Clear Yes/No Answer with Defense Prompt
Table 11: Model outputs missing Yes/No tokens with defense prompt.
@@
HealthCareMagic flan 0 4000 $0$ 4000 00.00%
llama 0 4000 00.00% 3868 4000 96.70%
@@
Enron flan 0 4000 $0$ 4000 00.00%
llama 0 4000 00.00% 3805 4000 95.12%
@@
The number of model outputs that did not contain either of the ”Yes” or ”No” tokens when employing a defense in the RAG template (using attack prompt #2) for each dataset and model combination are shown in Table 11. This is compared to the corresponding model outputs when using attack prompt #2 without the defense. The number of missing answers increases significantly for the llama and mistral models, and remains the same (0) for flan.